A line is passing through the points (-1,1) adn (-2,-1).
What is the lope and what is the equation of the line and what are the x and y intercepts?

Answer by checkley71(8403)

THE SLOPE OF THIS LINE IS (-1-1)/-2+1) OR -2/-1 OR A SLOPE OF 2
SUBSTITUTINS ONE SET OF POINTS WE GET THE Y INTERCEPT USING THE FORMULA
Y=mX+B OR 1=2(-1)+b OR 1=-2+b OR b=1+2 OR b=3
NOW SUBSTITUTE 0 FOR Y & WE GET 0=2X+3 OR 2X=-3 OR X=-3/2 OR X=-1.5
THUS WE HAVE A SLOPE OF (2) A Y INTERCEPT OF (3) & AN X INTERCEPT OF (-1.5)